# Benchmarking Mental State Representations in Language Models

### v1

@@ Abstract
While numerous works have assessed the generative performance of language models (LMs) on tasks requiring Theory of Mind reasoning, research into the models’ internal representation of mental states remains limited. Recent work has used probing to demonstrate that LMs can represent beliefs of themselves and others. However, these claims are accompanied by limited evaluation, making it difficult to assess how mental state representations are affected by model design and training choices. We report an extensive benchmark with various LM types with different model sizes, fine-tuning approaches, and prompt designs to study the robustness of mental state representations and memorisation issues within the probes. Our results show that the quality of models’ internal representations of the beliefs of others increases with model size and, more crucially, with fine-tuning. We are the first to study how prompt variations impact probing performance on theory of mind tasks. We demonstrate that models’ representations are sensitive to prompt variations, even when such variations should be beneficial. Finally, we complement previous activation editing experiments on Theory of Mind tasks and show that it is possible to improve models’ reasoning performance by steering their activations without the need to train any probe.
Machine Learning, ICML
@@ 1 Introduction
Modern language models (LMs) trained on next token prediction have demonstrated impressive capabilities, spanning coding, mathematical reasoning, fact verification, and embodied interaction (Wei et al., 2022; Bubeck et al., 2023). As these models are designed with the ultimate goal of collaborating with humans, it becomes imperative that they complement these skills with an understanding of humans, in particular their beliefs, emotions, desires, and intentions (Li et al., 2023a). Core to this understanding is Theory of Mind (ToM) – the ability to attribute mental states to oneself and others (Premack & Woodruff, 1978). ToM is essential for effective communication and cooperation with other agents, facilitating interaction and learning from feedback and demonstrations (Saha et al., 2023). Given its significance, ToM has emerged as a critical milestone in artificial intelligence (AI) and an important capability when evaluating cutting-edge LMs (Bubeck et al., 2023). Interest in LMs’ generative performance on tasks requiring ToM reasoning has resulted in a wide variety of benchmark datasets, typically involving question-answering tasks (Le et al., 2019; Gandhi et al., 2023; Kim et al., 2023; He et al., 2023; Tan et al., 2024; Xu et al., 2024).
Despite showing improved performance on ToM benchmarks compared to earlier models, modern LMs are still far from perfect (Sap et al., 2022). Text generated by LMs often contains errors that limit their performance on ToM tasks (Martindale et al., 2019). Previous work has shown that it is sometimes possible to still obtain correct predictions by probing LMs’ internal representations (Li et al., 2021; Liu et al., 2023b; Gurnee et al., 2023). In particular, Zhu et al. (2024) have shown that LMs, when prompted with a story and a belief statement, can represent beliefs from their own perspective and, to a lesser extent, from the perspective of a character in the story. Their work is an important first step towards understanding how LMs represent mental states, but it is limited in the number of models and settings studied, leaving many questions unanswered.
Building and extending on (Zhu et al., 2024), we benchmark mental state representations of self and others in language models through extensive experiments of different LM families, model sizes, and fine-tuning approaches. Specifically, we design a set of experiments to address the following research questions: RQ1. What is the relation between model size and probing accuracy? RQ2. Does fine-tuning with instruction-tuning (Wei et al., 2021) and/or reinforcement learning from human feedback (Christiano et al., 2017; Ouyang et al., 2022, RLHF) have an effect on probing accuracy? RQ3. Are models’ internal representations of beliefs sensitive to prompt variations? RQ4. Is there a risk of probes memorising training data due to the large dimensionality of LM representations? RQ5. Can we enhance LMs’ performance by editing their activations without training dedicated probes?
To answer RQ1, we perform probing experiments on two families of LMs, Llama-2 (Touvron et al., 2023), and Pythia (Biderman et al., 2023), ranging from models with 70 million to 70 billion parameters. To address RQ2, we compare the probing performance of models pre-trained solely on next token prediction with models that have been fine-tuned using instruction-tuning and/or RLHF. Our experiments reveal that probing accuracy on the beliefs of others increases with model size and, more crucially, with fine-tuning. To answer RQ3, we repeat our probing experiments using different variations of the prompt used by Zhu et al.. Specifically, we consider two variations that are expected to negatively impact LMs’ representations (Random, Misleading), and two that are supposed to have a positive influence (Time Specification, Initial Belief). By conducting these experiments, our work is the first to explore the sensitivity of LMs’ representations to prompting in the context of ToM. Our findings demonstrate that models’ representations are sensitive to prompt variations, even when such variations should be beneficial. To address RQ4, we compare our trained probes with a second set of probes trained only on the representations’ first top $k$ principal components. This requires learning much fewer parameters and eliminates any possible memorisation issue. We find no strong evidence of memorisation in the probes, as it is possible to recover most of the accuracy by training probes on a small subset of principal components of models’ representations. We formulate RQ5 as a follow-up question to Zhu et al. (2024) who found that probes trained to predict beliefs can be used to steer models’ activation using inference-time intervention (Li et al., 2023c, ITI) to improve LMs’ downstream performance on ToM tasks. In contrast, we show that by using contrastive activation addition (Rimsky et al., 2023, CAA), we can steer models’ activations without the need to train any probe and, in a more generalisable way, obtain significant performance improvements across different ToM tasks.
@@
    We report extensive probing experiments with various types of LMs with different model sizes and fine-tuning approaches, showing that the quality of models’ internal representations of the beliefs of others increases with model size and, more crucially, fine-tuning.
@@
    We are the first to study how prompt variations impact belief probing performance, showing that models’ representations are sensitive to prompt variations, even when such variations should be beneficial.
  3. 3.
    We show that by using contrastive activation addition it is possible to improve models’ reasoning performance by steering their activations without the need to train any probe.
@@ 2 Related Work
Theory of Mind (ToM) has been studied in cognitive science and psychology for decades (Gurney et al., 2021). Mirroring efforts to understand ToM in humans, an increasing number of works in the computational sciences have investigated means to equip AI with similar capabilities. Previously proposed models that aim to implement a machine ToM have been based on partially observable Markov decision processes (POMDP) (Doshi et al., 2010; Han & Gmytrasiewicz, 2018), Bayesian methods (Baker et al., 2011, 2017) and deep learning methods (Rabinowitz et al., 2018; Bara et al., 2021; Wang et al., 2022; Duan et al., 2022; Liu et al., 2023a; Bortoletto et al., 2024b, a). Recent advances in LMs have sparked interest in evaluating their ToM capabilities. Various benchmarks have been proposed, aiming to measure LMs’ ability to understand and reason about the beliefs, goals, and intentions of others (Le et al., 2019; He et al., 2023; Kim et al., 2023; Gandhi et al., 2023; Xu et al., 2024; Tan et al., 2024). Additionally, efforts have been made to enhance LMs’ ToM through prompting techniques (Zhou et al., 2023b; Moghaddam & Honey, 2023; Wilf et al., 2023). A new direction of research explores LMs’ internal representation of mental states. Zhu et al. (2024) demonstrated that LMs linearly encode beliefs from different agents’ perspectives, and manipulating these representations can enhance ToM task performance. While Zhu et al.’s work is a crucial initial step, our work dives deeper into LMs’ internal belief representations, offering a broader insight into these mechanisms.
@@
Initially proposed by Alain & Bengio (2017), probing has emerged as a common method for determining if models represent particular features or concepts. In the realm of LMs, numerous works used probing to demonstrate that these models acquire rich linguistic representations. These representations span syntactic and semantic concepts such as syntactic categories, dependency relations, co-reference, and word meaning (Conneau et al., 2018; Tenney et al., 2018, 2019; Rogers et al., 2021; Li et al., 2021; Hernandez & Andreas, 2021; Marks & Tegmark, 2023; Liu et al., 2023b). A separate line of work explored if and how LMs represent the world, i.e., whether they possess a world model. Li et al. (2021) showed that LMs track the states of entities within a context. Other works showed that LMs exhibit representations reflecting non-linguistic concepts in the world, which LMs have never observed (Abdou et al., 2021; Patel & Pavlick, 2022; Li et al., 2023b; Nanda et al., 2023). An emergent line of work that is particularly relevant to our work used probing to explore if LMs have agent models, for example, if they can represent beliefs of self and others (Zhu et al., 2024; Bortoletto et al., 2024a). While representing an important first step towards understanding the internals of Theory of Mind in LMs, experiments in (Zhu et al., 2024) are limited in settings and models considered. In this work, we contribute with extensive experiments that employ a wider variety of LMs and a wider range of settings. Furthermore, we also explore possible memorisation issues in the probes.
@@
Research on prompt robustness in LMs is still in its infancy but has quickly sparked much interest. On one hand, previous work has shown that LMs are vulnerable to prompt alterations like token deletion or reordering (Ishibashi et al., 2023), biased or toxic prompts (Shaikh et al., 2023) and similarity to training data (Razeghi et al., 2022). On the other hand, instruction-tuned models have proved to be more robust against prompt variation, even when using misleading instructions (Webson & Pavlick, 2022). Other works have shown the importance of input-output format (Min et al., 2022) and of demonstration example ordering for few-shot performance (Zhao et al., 2021; Lu et al., 2022; Zhou et al., 2023a). In this work, we shift our focus from analysing how sensitive model outputs are to how model representations change. Our work, along with (Gurnee et al., 2023), is one of the first to explore how prompt design affects how accurately models represent concepts. In particular, Gurnee et al. (2023) have studied whether LMs’ representations of space and time are robust to prompt variations. In stark contrast, we explore for the first time the effect of prompt variations on how models represent mental states internally.
@@
Recent advancements in NLP have introduced innovative techniques for controlling and manipulating text generation models. While weight editing proposed to modify models’ weights (Meng et al., 2022; Ilharco et al., 2022; Orgad et al., 2023), activation editing has emerged as an alternative way to influence model behaviour without any additional fine-tuning (Li et al., 2023b; Hernandez et al., 2023). This approach involves manipulating the internal representations of models to direct their outputs towards desired outcomes. One notable method in this domain is inference-time intervention (Li et al., 2023c, ITI), which has been proposed to enhance truthfulness in LMs. ITI involves training linear probes on contrastive question-answering datasets to identify “truthful” attention heads and then shifting attention head activations during inference along the identified truthful directions. In contrast, activation addition (Turner et al., 2023, AA) and contrastive activation addition (Rimsky et al., 2023, CAA) offer ways to generate steering vectors by only using LMs’ intermediate activations. Zhu et al. have used ITI to show that it is possible to manipulate LMs’ internal representations of mental states (Zhu et al., 2024). In this work, we show that using CAA can further improve LMs’ ToM capabilities without the necessity of training any probe. Remarkably, CAA operates at the residual stream level, eliminating the need for a fine-grained search over attention heads.
@@ 3.1 Probing
Figure 1: Example of false belief from our probing datasets. The labels $z_{p}$ and $z_{o}$ correspond to $\mathcal{D}_{p}^{P}$ and $\mathcal{D}_{o}^{P}$, respectively. By manipulating the protagonist’s percepts after the causal event we obtain two scenarios: true belief and false belief.
In line with previous work (Zhu et al., 2024) we linearly decode belief status from the perspective of different agents by using probing (Alain & Bengio, 2017). Probing involves localising specific concepts in a neural model by training a simple classifier (called a probe) on model activations to predict a target label associated with the input data. To provide a formal definition, we adopt a similar notation to the one introduced in (Belinkov, 2022). Let us define an original model $f:x\mapsto\hat{y}$ that is trained on a dataset $\mathcal{D}^{O}=\{x^{(i)},y^{(i)}\}$ to map input $x$ to output $\hat{y}$. Model performance is evaluated by some measure, denoted $\textsc{Perf}(f,\mathcal{D}^{O})$. A probe $g:f_{l}(x)\mapsto\hat{z}$ maps intermediate representations of $x$ in $f$ at layer $l$ to some property $\hat{z}$, which is the label of interest. The probe $g$ is trained on a probing dataset $\mathcal{D}^{P}=\{x^{(i)},z^{(i)}\}$ and evaluated using some performance measure $\textsc{Perf}(g,f,\mathcal{D}^{O},\mathcal{D}^{P})$. In our case, $f$ is an autoregressive language model that given a sequence of tokens $x$ outputs a probability distribution over the token vocabulary to predict the next token in the sequence. Our probe is a logistic regression model $\hat{z}=\sigma(Wa_{l}+b)$ trained on neural activations $f_{l}(x)=a_{l}$ to predict belief labels $y=\{0,1\}$.
@@ 3.2 Dataset
Following Zhu et al. (2024) we use the BigToM benchmark (Gandhi et al., 2023). BigToM is constructed using GPT-4 (Achiam et al., 2023) to populate causal templates and combine elements from these templates. Each causal template is set up with a context and a description of the protagonist (e.g. “Noor is working as a barista […]”), a desire (“Noor wants to make a cappuccino”), a percept (“Noor grabs a milk pitcher and fills it with oat milk”), and a belief (“Noor believes that the pitcher contains oat milk”). The state of the world is changed by a causal event (“A coworker swaps the oat milk in the pitcher with almond milk”). The dataset constructs different conditions by changing the percepts of the protagonist after the causal event, which will result in different beliefs. In this work, we focus on the Forward Belief setting proposed by (Zhu et al., 2024) in which models have to infer the belief of the protagonist given the percepts of the causal event, $P(\mathrm{belief}|\mathrm{percepts})$. We report additional details in Section A.1.1
@@
We consider two probing datasets: $\mathcal{D}_{p}^{P}=\{x^{(i)}_{p},z^{(i)}_{p}\}$, where the labels $z^{(i)}_{p}$ correspond to ground-truth beliefs from the protagonist perspective, and $\mathcal{D}_{o}^{P}=\{x^{(i)}_{o},z^{(i)}_{o}\}$, where the labels $z^{(i)}_{o}$ reflect the perspective of an omniscient oracle. $\mathcal{D}_{p}^{P}$ and $\mathcal{D}_{o}^{P}$ are built by pairing each story in BigToM with a belief statement, as shown in Figure 1. After prompting the model with a story-belief pair we cache the residual stream activations at the final token position for all residual streams (Figure 5).

### v3

# Brittle Minds, Fixable Activations: Understanding Belief Representations in Language Models

@@ Abstract
Despite growing interest in Theory of Mind (ToM) tasks for evaluating language models (LMs), little is known about how LMs internally represent mental states of self and others. Understanding these internal mechanisms is critical – not only to move beyond surface-level performance, but also for model alignment and safety, where subtle misattributions of mental states may go undetected in generated outputs. In this work, we present the first systematic investigation of belief representations in LMs by probing models across different scales, training regimens, and prompts – using control tasks to rule out confounds. Our experiments provide evidence that both model size and fine-tuning substantially improve LMs’ internal representations of others’ beliefs, which are structured – not mere by-products of spurious correlations – yet brittle to prompt variations. Crucially, we show that these representations can be strengthened: targeted edits to model activations can correct wrong ToM inferences.
Brittle Minds, Fixable Activations:
Understanding Belief Representations in Language Models
Matteo Bortoletto  Constantin Ruhdorfer  Lei Shi  Andreas Bulling University of Stuttgart, Germany matteo.bortoletto@vis.uni-stuttgart.de
@@ 1 Introduction
Language models (LMs) trained on next token prediction have demonstrated impressive capabilities across various tasks, spanning coding, math, and embodied interaction  (Wei et al., 2022; Bubeck et al., 2023). As these models are designed with the ultimate goal of collaborating with humans, it becomes imperative that they complement these skills with an understanding of humans. Core to this understanding is Theory of Mind (ToM) – the ability to attribute mental states to oneself and others (Premack and Woodruff, 1978). ToM is essential for effective communication and cooperation with other agents, facilitating interaction and learning from feedback and demonstrations (Saha et al., 2023). Given its significance, computational ToM has emerged as a key capability when evaluating cutting-edge LMs (Ma et al., 2023; Shapira et al., 2024; Chen et al., 2025).
Figure 1: ToM tasks are challenging for LMs, but correct predictions can sometimes be recovered by probing their internal representations. We study how internal representations of beliefs of self and others emerge in 12 LMs, and show that these representations are structured yet brittle to prompts, and can be strengthened with a steering vector to fix incorrect ToM inferences.
Despite the improved performance on ToM benchmarks compared to earlier models, modern LMs are still far from perfect (Sap et al., 2022). Text generated by LMs often contains errors that limit their performance on ToM tasks. Zhu et al. (2024) showed that probing LMs’ internal representations can sometimes recover correct belief inferences, with models like Mistral-7B-Instruct Jiang et al. (2023) and DeepSeek-7B-Chat Bi et al. (2024) capturing beliefs from both their own and others’ perspectives. While promising, this remains a preliminary step: it examines only single-sized, fine-tuned models, leaves possible confounds uncontrolled, and ignores how subtle changes in prompting affect belief representations. As a result, we still lack a clear understanding of how internal belief representations differ across models, whether they reflect true ToM or spurious patterns, and how robust they are to prompts.
To address these gaps, we pose four key research questions and present evidence for each. We begin by studying how emergence scales across models:
RQ1
Do internal belief representations emerge similarly in different LMs, and are they affected by model size and training regime?
Finding training regimes or scales that are more conducive to belief reasoning can guide future model development toward more reliable ToM behaviour. However, it is also crucial to verify if representations are structured, indicating genuine modelling of mental states, or spurious:
RQ2
Are LMs’ internal belief representations structured or the result of spurious correlations?
This distinction is essential for determining if representations reflect a genuine understanding of beliefs or only exploit statistical patterns that happen to correlate with correct answers in the training data. This is also crucial for alignment and safety, as misaligned mental state attributions may not appear overtly in text – leading to false signals of understanding. Equally important is that models can maintain robust belief attributions:
RQ3
Are LMs’ internal belief representations robust?
Fragile representations may break under slight variations, leading to inconsistent or unsafe behaviour in real-world applications involving social reasoning or user interaction. Strengthening these representations, then, offers a promising path toward improving their reliability:
RQ4
Can we strengthen LMs’ internal belief representations to improve their performance?
To answer these research questions, we perform probing and activation editing experiments using 12 LMs (Figure 1). We first compare base models with those fine-tuned via SFT and/or RLHF Ouyang et al. (2022)(RQ1), finding that belief representations emerge in consistent patterns across models, improve with model size, and – especially in smaller models – benefit significantly from fine-tuning. To provide evidence that LMs’ belief representations are structured (RQ2), we show that (1) probes trained on randomly permuted labels perform at chance – confirming selectivity, and (2) probes trained on top-$k$ principal components still recover most accuracy for $k\ll d_{model}$. Next, we test robustness (RQ3) using varied prompts. Surprisingly, semantically neutral changes can reduce accuracy, revealing that representations of others’ beliefs are brittle to prompts. However, we show that it is possible to strengthen models’ representation by using contrastive activation addition (Rimsky et al., 2023, CAA), obtaining significant performance improvements across different ToM tasks (RQ4).
@@
    We provide extensive probing experiments across 12 LMs, suggesting that representations of others’ beliefs improve with size and fine-tuning, and that these representations are structured yet brittle to prompt variations.
@@
    We show that we can strengthen models’ representations by using contrastive activation addition and improve their ToM performance.
@@ 2 Related Work
Theory of mind has been studied in AI for more than a decade (Baker et al., 2009; Rabinowitz et al., 2018; Bara et al., 2021; Bortoletto et al., 2024a, b, c). Various benchmarks have been proposed, aiming to measure LMs’ ability to understand and reason about the beliefs, goals, and intentions of others (Le et al., 2019; He et al., 2023; Kim et al., 2023; Gandhi et al., 2023; Xu et al., 2024; Tan et al., 2024; Sclar et al., 2023; Ma et al., 2023; Wu et al., 2023). Additionally, efforts have been made to enhance LMs’ ToM through prompting techniques (Zhou et al., 2023b; Moghaddam and Honey, 2023; Wilf et al., 2023). Our work dives deeper into LMs’ internal belief representations, offering a broader insight into these mechanisms that go beyond surface-level performance.
@@
Initially proposed by Alain and Bengio (2017), probing is a widely used method for determining if models represent particular features or concepts. In the realm of LMs, numerous works used probing to demonstrate that these models acquire rich linguistic representations – spanning semantic concepts such as syntactic categories, dependency relations, co-reference, and word meaning (Conneau et al., 2018; Tenney et al., 2018, 2019; Rogers et al., 2021; Li et al., 2021; Hernandez and Andreas, 2021; Marks and Tegmark, 2023; Liu et al., 2023). A separate line of work explored if LMs possess a world model (Li et al., 2021; Abdou et al., 2021; Patel and Pavlick, 2022; Li et al., 2023a; Nanda et al., 2023). An emergent line of work that is relevant to our work used probing to explore if LMs have agent models, for example, if they can represent beliefs of self and others (Zhu et al., 2024; Bortoletto et al., 2024a). In this work, we contribute with extensive experiments that characterise models’ representations of beliefs along different axes: emergence, structure, robustness, and steerability.
@@
Previous work has shown that LMs are vulnerable to prompt alterations like token deletion or reordering (Ishibashi et al., 2023), biased or toxic prompts (Shaikh et al., 2023) and similarity to training data (Razeghi et al., 2022). Other works have shown the importance of input-output format (Min et al., 2022) and of demonstration example ordering for few-shot performance (Zhao et al., 2021; Lu et al., 2022; Zhou et al., 2023a). In this work, we shift our focus from analysing how sensitive model outputs are to how model representations change Gurnee and Tegmark (2024). In particular, we explore for the first time the effect of prompt variations on how models internally represent mental states.
@@
Activation editing has emerged as a way to influence model behaviour without any additional fine-tuning (Li et al., 2023a; Hernandez et al., 2023). One notable method in this domain is inference-time intervention (Li et al., 2023b, ITI), which involves training linear probes on contrastive question-answering datasets to identify “truthful” attention heads and then shifting their activations during inference along the identified truthful directions. In contrast, activation addition (Turner et al., 2023, AA) and contrastive activation addition (Rimsky et al., 2023, CAA) generate steering vectors by only using LMs’ activations. Zhu et al. has used ITI to show that it is possible to manipulate LMs’ internal representations of mental states. In this work, we show that using CAA can further improve LMs’ ToM capabilities while eliminating the need for a fine-grained search over attention heads.
@@ 3.1 Probing
We linearly decode belief status from the perspective of different agents by using probing (Alain and Bengio, 2017). Probing involves localising specific concepts in a neural model by training a simple classifier (called a probe) on model activations to predict a target label associated with the input data. To provide a formal definition, we adopt a similar notation to the one introduced in (Belinkov, 2022). Consider an original model $f:x\mapsto\hat{y}$ that is trained on a dataset $\mathcal{D}^{O}=\{x^{(i)},y^{(i)}\}$ to map input $x$ to output $\hat{y}$. Model performance is evaluated by some measure, denoted $\textsc{Perf}(f,\mathcal{D}^{O})$. A probe $g_{l}:f_{l}(x)\mapsto\hat{z}$ maps intermediate representations of $x$ in $f$ at layer $l$ to some property $\hat{z}$, which is the label of interest. The probe $g_{l}$ is trained on a probing dataset $\mathcal{D}^{P}=\{x^{(i)},z^{(i)}\}$ and evaluated using some performance measure $\textsc{Perf}(g_{l},f,\mathcal{D}^{O},\mathcal{D}^{P})$. In our case, $f$ is an autoregressive language model that, given a sequence of tokens $x$, outputs a probability distribution over the token vocabulary to predict the next token in the sequence. Our probe is a logistic regression model $g_{l}:\hat{z}=Wa_{l}+b$ trained on neural activations $f_{l}(x)=a_{l}$ to predict binary belief labels $y=\{0,1\}$.
@@ 3.2 Dataset
We use BigToM (Gandhi et al., 2023), a question-answering dataset constructed by populating causal templates and combining elements from these templates. Each causal template is set up with a context and a description of the protagonist (e.g. “Noor is working as a barista […]”, see Story in Figure 2), a desire (“Noor wants to make a cappuccino”), a percept (“Noor grabs a milk pitcher and fills it with oat milk”), and a belief (“Noor believes that the pitcher contains oat milk”). The state of the world is changed by a causal event (“A coworker swaps the oat milk in the pitcher with almond milk”). The dataset constructs different conditions by changing the percepts of the protagonist after the causal event, which will result in different beliefs. Similar to (Zhu et al., 2024), we focus on the Forward Belief setting in which models have to infer the belief of the protagonist given the percepts of the causal event, $P(\mathrm{belief}|\mathrm{percepts})$. We report additional details in Section A.1.1
Figure 2: Example of false belief from our probing datasets. The labels $z_{p}$ and $z_{o}$ correspond to $\mathcal{D}_{p}^{P}$ and $\mathcal{D}_{o}^{P}$, respectively. By manipulating the protagonist’s percepts after the causal event, we obtain two scenarios: true belief and false belief.
@@
We consider two probing datasets: $\mathcal{D}_{p}^{P}=\{x^{(i)}_{p},z^{(i)}_{p}\}$, where the labels $z^{(i)}_{p}$ correspond to ground-truth beliefs from the protagonist perspective, and $\mathcal{D}_{o}^{P}=\{x^{(i)}_{o},z^{(i)}_{o}\}$, where the labels $z^{(i)}_{o}$ reflect the perspective of an omniscient oracle. $\mathcal{D}_{p}^{P}$ and $\mathcal{D}_{o}^{P}$ are built by pairing each story in BigToM with a belief statement, as shown in Figure 2. After prompting the model with a story-belief pair $x$ we cache the residual stream activations $f_{l}(x)$ at the final token position for all residual streams (see Figure 6).